OpenStudy (imalexlowe):

If f(x)=2(x)^2+5√(x+2), complete the following statement: f(2)=

sam (.sam.):

\[If ~~~~~~~~~~f(x)=2(x)^2+5\sqrt{(x+2)}\] \[Then ~~~~~~~~~~f(2)=2(2)^2+5\sqrt{(2+2)}\]